12. Determine if the functions are even, odd, or neither. Show your work to justify your (a) f(x) = x4 + 2x2-1 (b) f(x)+2
Answer #1

A) f(x)=x^4+2x^2-1

f(-x) can be obtained by replacing x by -x in the equation for f(x).

If we find out that f(x) = - f(x), then the function is odd.

If f(x) = f(- x), then the function is even.

Here

f(-x)=(-x)^4+2(-x)^2-1

f(-x)=x^4+2x^2-1

Thus we find out that  f(-x)=f(x).

So the given function is even function.

B) f(x)=x^{-1} +2

f(x)=\frac {1}{x}+2

We need to get f(- x).

f(-x)=\frac {1}{-x}+2

f(-x)=-\frac {1}{x}+2

We find out that

f(-x)\neq f(x) \, and\, also \,f(-x) \neq-f(x)

Thus the function is neither odd nor nor even.